2. The historical demand for a product is below Plese forcast the drmapr 300 (fid the company Month Profit in Year 20167 Sales (Use a 3-month simple moving z 23 (bUse a3month weighted moving average w 3 (March) March forecast 29 Find trend line and then foecast demand O p
Answer #1

a) Simple moving average= Sum of n Sales / n.

So (23+27+31) / 3 = 27

b) Weighted moving average: (weight %1 X Sales1) + (weight %2 X Sales2) + (weight %3 X Sales3)
Hence (0.6 X 23) + (0.3 X 27) + (0.1 X 31) = 24

c) March Actual Sale = 31 , a=0.8 ,March Forecast= 29

April Forecast = March forecast + a X (Actual Sales - March Forecast)
   = 29 + 0.8 (31-29)
   = 30.6

d) information missing. required value of Beta and trend of January